OpenStudy (anonymous):

3. A car speeds up from 10.0 m/s to 45.0 m/s in 4.0 s. What is its acceleration? In your response, show your work, the given and unknown quantities, the equation, and any necessary algebra. Express your answer with the correct number of significant digits

OpenStudy (anonymous):

@ganeshie8

OpenStudy (shamim):

Required equation is a=(v-u)/t

OpenStudy (shamim):

Here a=accekeration v=final velocity u=initial velocity t=time

OpenStudy (shamim):

Given u=10m/s v=45m/s t=4sec a=?

OpenStudy (shamim):

Response plz!!!!!